# Connection between gravity and quantum mechanics: an algebraic approach to cosmological issues

### v1

@@ III Planck equivalence and corresponding relations
It is well-recognized that a hallmark of any well-established theory is its natural derivation of fundamental constants, equivalences, and relationships. In our work, we independently developed a significant system of Planck units, necessitating a thorough examination of the physical implications and extensive consequences that arise from it. For instance, this approach reveals a series of astonishing formulas and equivalences demonstrating the system’s inherent depth and coherence. The formulas and equivalences are:
@@
IV Energy and entropy radiation of black holes
The description of a black hole closely aligns with the renowned “membrane formalism” proposed by Price and Thorne [39], which reconceptualizes the dynamics of the black hole horizon as a two-dimensional “stretched horizon” that operates as a physical membrane within three-dimensional space. In our case, we identify two significant differences from that perspective [4].
@@ IV Energy and entropy radiation of black holes
2. In our framework, the significant gravitational squeezing force, which explained as below:
@@
which is precisely balanced by analogous repulsive quantum-mechanical counterpart as:
@@
This has profound implications for forming a stable quantum gravitational state that prevents singularities. An equilibrium between gravitational compression and quantum repulsion can establish conditions that preserve quantum information at the horizon, consistent with concepts from holography and the AdS/CFT correspondence [52, 7, 53, 54]. The Holographic principle, proposed by Gerard ’t Hooft and further developed by Leonard Susskind, proposes that all information contained within a volume can be encoded on its boundary, analogous to a three-dimensional hologram represented on a two-dimensional surface. The AdS/CFT correspondence, introduced by Juan Maldacena in 1999, exemplifies this principle’s application. This work shows how two different theories of gravity work together in AdS/CFT at its boundary. It also makes a link between quantum field theory and gravity, which makes it easier to study quantum gravity, black holes, and condensed matter physics.
However, this remarkable equilibrium is periodically interrupted by spontaneous quantum transitions, as described by the work of Bekenstein and Mukhanov [21]. Let’s delve deeper to understand what this means in the context of the structure formation of the universe. According to our understanding, quantum fluctuations are critical in establishing the initial density variations in the early universe that eventually grow into galaxies and large-scale structures. These fluctuations create a “spontaneous transition” in the nascent universe by introducing new phases or areas with slightly differing energy densities. A “spontaneous quantum transition” could occur when the universe’s energy density reaches a critical threshold, causing it to “jump” to a new state. This could be analogous to vacuum decay or phase transitions in early cosmology, in which spacetime “selects” regions with lower energy, resulting in the expansion or reconfiguration of spacetime domains. These transitions occur from an initial $n_{s}$ energy level to a lower $n_{s}-1$ state, resembling electron transitions in the hydrogen atom. The conservation of momentum drives such transitions, which involve the emission of two particles. Gravitational radiation ($\Gamma_{s}$), which appears in quadrupolar and octupolar wave modes, also coincides with these quantum leaps ($\gamma_{s}$). The fundamental framework governing these transitions is outlined as follows [4]:
@@
$\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)=\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)+\Delta E_{s,s-1}\leavevmode\nobreak\ gr(rad)$
where $qu(rad)$ and $gr(rad)$ represents quantum radiation and gravitational radiation respectively. Also, note that throughout the article, ($qu$), ($gr$), and ($rad$) represent “quantum”, “gravitational”, and “radiation” respectively. Let us evaluate the following terms:
@@
and from Eq. (14), we get
$\Delta E_{s,s-1}\leavevmode\nobreak\ gr(rad)=\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)-\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)$
@@
$=\Delta E_{s,s-1}\leavevmode\nobreak\ (qu)\big{(}\frac{1}{4n_{s}}+\frac{1}{8n_{s}^{2}}+...\big{)}$ (16)
@@
$\Delta S_{s,s-1}\leavevmode\nobreak\ qu(rad)=k_{s}2\leavevmode\nobreak\ (qubit);$
$\Delta S_{s,s-1}\leavevmode\nobreak\ gr(rad)=k_{s}2\leavevmode\nobreak\ (qubit)\big{(}\frac{1}{4n_{s}}+\frac{1}{8n_{s}^{2}}+...\big{)},$
@@
    Considering the straightforward physical expression of the acceleration $a_{s}$, linked to the dynamical process of a black hole’s evaporation
@@
    Using the previously stated equivalence $T_{p}=\frac{a_{p}}{ck_{B}}\hbar$, we can derive the remarkable expression that relates the temperature and acceleration as:
    $T_{s}(rad)=\frac{E_{p}}{4ck_{B}\sqrt{n_{s}}}$
@@ V PBHs Lifetime and Luminosity: Addressing the Information Loss Paradox
PBHs provide a distinctive platform to test theories of quantum gravity and black hole evaporation. Here, we aim to establish a connection between general relativity and quantum physics by examining lifetimes, luminosities, and the potential elimination of information-loss paradoxes. The time of transition of PBHs from one quantum state to the underlying one can be calculated using the strong conditions of causality and spherical geometry. It is equal to half the length of the geodesic line divided by the speed of light [4],
$\Delta t_{s}=\frac{\pi R_{s}}{c}=2\pi t_{p}\sqrt{n_{s}}.$ (22)
The aforementioned expression for black hole lifetime can be readily found using this formula as:
$\tau_{s}=\sum_{n=1}^{n_{s}}2\pi t_{p}\sqrt{n}$
@@
where the usual lifetime and luminosity of the black hole can be expressed as: $\tau(Hawking)\simeq\frac{5120\pi G^{2}}{\hbar c^{4}}M^{3}$ and $L(Hawking)\simeq\frac{\hbar c^{6}}{15360\pi G^{2}M^{2}}$ [14, 15, 57]. This model suggests that black holes undergoing quantum transitions differ substantially from Hawking’s semi-classical evaporation model. Quantum processes result in more intense emissions than continuous black-body radiation, reducing the black hole’s lifetime by a factor of 3840 and increasing its luminosity by the same factor due to discrete energy jumps. Consequently, without external energy input (e.g., through accretion) to offset this intense quantum emission, the black hole undergoes rapid evaporation. This finding challenges Hawking’s assumption that black holes radiate as black bodies under the Stefan-Boltzmann law, showing instead that quantum transitions produce significantly higher radiation intensities.
Now we calculate the energy of each emitted particle:
@@
From the above Eq. (25), one can easily find that their emitted wavelength is much more than $R_{s}$ :
@@
Because the wavelength of the emitted particles is greater than the Schwarzschild radius, the black hole’s strong gravitational field does not prevent the particles from tunneling out. In our scenario, each quantum transition results in the emission of two particles, each containing a discrete amount of energy and information. As a result, each emission causes the black hole to lose two bits of quantum information. However, in conventional black hole evaporation, Hawking radiation is merely thermal and contains no information about the black hole’s internal state, resulting in the so-called information loss paradox. In contrast, our model may encode information in the emitted particles, allowing information about the black hole’s interior states to escape with each two-particle emission. As the black hole emits radiation, it loses mass, energy, and information in a quantised manner, resolving the information loss dilemma by allowing information to leave the black hole in the form of emitted quantum bits (qubits).
@@ VI Quantum Absorption of Free Light: Origins of Cosmic Disks and Jets
Considering the quantum characteristics of condensed light in black holes, it is evident that only free light photons with a wavelength not exceeding the geodesic length of the black hole are subject to accretion by quantum absorption. In other words, black holes consume only free light or condensed light (such as other black holes through collisions and mergers). All ordinary matter is melted in accretion disks and expelled in jets, accompanied by swirling magnetic fields through “funnel” and “slingshot” effects. In this scenario, the wavelength of an absorbed photon cannot exceed the geodesic length; we get a condition on wavelength:
$\lambda_{\gamma}\leavevmode\nobreak\ (absorp)\leq 2\pi R_{s}=\lambda_{sm}$ (27)
@@
If, during the specified time interval $\Delta t_{s}$ (22), which corresponds to the duration of the quantum transition from energy level $n_{s}$ to $n_{s-1}$, the black hole does not absorb the aforementioned free light quantum of energy, it will continue to evaporate. If it receives, it grows in accordance with the law:
@@
It means when the temperature of the universe ($T_{\gamma}$) falls below twice the temperature of a black hole, it commences evaporation. Furthermore, it is essential to comprehend how, in the nascent universe, the demise of one generation of PBHs facilitated the emergence of the subsequent generation [14, 15, 58, 59]. We must adhere to these paramount cosmological criteria in our explanation of the formation of contemporary dark matter, which comprises black holes of asteroid-like mass [60]. Another significant consequence of these interactions is that an evaporating black hole can transfer its electromagnetic radiation energy to another black hole only if the mass of the latter is double that of the former [61],
@@
These correlations elucidate how, in the primordial phases of the universe’s growth, there occurred a cyclic transfer of energy from a dying, more abundant yet lighter generation of black holes to a new, somewhat less abundant but heavier generation. This is essential for comprehending both the primordial formation of dark matter’s structure and the initial “clumping” of the universe, as these consistent cycles of dark matter generation coincided with the hyper-exponential proliferation of future SMBHs (SMBHs), which served as the centers for the formation of subsequent galaxies.
VII Quantum Information Conservation and Gravitational Radiation from Black Hole collisions
We find that two-particle emission outside the black hole and accompanied by gravitational radiation can potentially resolve the ‘information loss paradox.’ John Preskill [18] and Leonard Susskind [62] define each emitted particle as a carrier of one bit of quantum information, in accordance with the principles of Shannon entropy [63]. This information is represented in two possible spin states, $|0>$ and $|1>$, as per the notation established by Paul Dirac [64]. This mechanism helps keep information safe and shows that the unitarity principle in quantum theory cannot be broken. This means that no information is lost during the process of black holes evaporating.
According to the conservation information law for a binary black hole system ($BH_{1}$,$BH_{2}$), which is derived from the basic unitarity principle of quantum theory, we can write:
@@ VII Quantum Information Conservation and Gravitational Radiation from Black Hole collisions
We derive an interesting formula for the energy of gravitational radiation resulting from the collision and merging of a binary black hole system,
@@
This formulation, expressed in terms of quantum numbers, succinctly corroborates Hawking’s esteemed area theorem [65] concerning colliding and merging black holes. During the merger, the resultant black hole experiences a ”ringdown” phase, during which it emits gravitational waves while stabilizing into a final state. Similar to a struck bell that resonates and then diminishes into quiet, the black hole emits energy as it reaches equilibrium—thus the evocative name “ringdown.”
@@
Now, we obtain the general expression for total gravitational radiation energy for N consecutive collisions and mergers of N black holes as :
@@
    When the mass of a SMBH at the center of a galaxy greatly exceeds the asteroid-like mass of PBHs situated in the SMBH’s “corona,” the dynamics of their mergers yield an interesting outcome. In these instances, nearly all the mass and energy of the tiny PBHs is converted into gravitational wave energy during the merger phase. This process demonstrates a significant transformation in which dark matter is permanently converted into dark energy within the universe. This process serves as a substantial validation of P. W. Anderson’s conjecture, highlighting the intricate interactions among different forms of matter and energy in the universe [67]. Moreover, as will be shown subsequently, these processes account for the universe’s present phase of accelerated expansion.
@@
    In contrast, Eqs. (36) and (37) suggest that the formation of a SMBH is unlikely during multiple collisions and mergers of black holes with nearly equal masses, as most of their total energy would be dissipated as gravitational radiation. It can be inferred that SMBHs primarily originated from the significant light energy associated with radiative processes in the early universe. We will explain below that these critical insights are vital for precisely characterising and comprehending the substantial role of PBHs and their gravitational radiation in cosmogenesis.
VIII Planck-scale physics, the hot Big Bang, and the Universe’s accelerating expansion

### v2

@@ III Planck equivalence and corresponding relations
It is well-recognized that a hallmark of any well-established theory lies in its natural derivation of fundamental constants, equivalences, and relationships. In our work, we independently developed a significant system of Planck units, necessitating a thorough examination of the physical implications and extensive consequences that arise from it. For instance, this approach reveals a series of astonishing formulas and equivalences demonstrating the system’s inherent depth and coherence. The formulas and equivalences are:
@@
IV Quantum emission of PBH: Energy and Entropy Radiation
The description of a black hole closely aligns with the renowned “membrane formalism” proposed by Price and Thorne [32], which reconceptualizes the dynamics of the black hole horizon as a two-dimensional “stretched horizon” that operates as a physical membrane within three-dimensional space. In our study, we identify two significant differences from that perspective, which warrant careful elucidation [44].
@@ IV Quantum emission of PBH: Energy and Entropy Radiation
2. Secondly, in our framework, the gravitational squeezing force, expressed as:
@@
is precisely balanced by analogous repulsive quantum-mechanical counterpart:
@@
This has profound implications for forming a quasi-stable quantum gravitational state that prevents singularities. An equilibrium between gravitational compression and quantum repulsion, analogous to the effect we found in white dwarf stars where electron degeneracy pressure holds the star against the gravitational collapse, can establish conditions that preserve quantum information at the horizon, consistent with concepts from holography and the AdS/CFT correspondence [45, 15, 46, 47]. The Holographic principle, proposed by Gerard ’t Hooft and further developed by Leonard Susskind, proposes that all information contained within a volume can be encoded on its boundary, analogous to a three-dimensional hologram represented on a two-dimensional surface. The AdS/CFT correspondence, introduced by Juan Maldacena in 1999, exemplifies this principle’s application. This work helps us to establish a link between quantum field theory and gravity, which makes it easier to study quantum gravity, black holes, and condensed matter physics.
However, this remarkable equilibrium is periodically interrupted by spontaneous quantum transitions, as described by the work of Bekenstein and Mukhanov [48]. This periodic interruption has a profound implication in the context of the structure formation of the universe. According to our understanding, quantum fluctuations are critical in establishing the initial density variations in the early universe that eventually grow into galaxies and large-scale structures. These fluctuations create a “spontaneous transition” in the nascent universe by introducing new phases or areas with slightly differing energy densities. A “spontaneous quantum transition” could occur when the universe’s energy density reaches a critical threshold, causing it to “jump” to a new state. This could be analogous to vacuum decay or phase transitions in early cosmology, in which spacetime “selects” regions with lower energy, resulting in the expansion or reconfiguration of spacetime domains. These transitions occur from an initial $n_{s}$ energy level to a lower $n_{s}-1$ state, resembling electron transitions in the hydrogen atom. The conservation of momentum drives such transitions, which involve the emission of two particles ($\gamma_{s}$) accompanied by gravitational radiation ($\Gamma_{s}$), in quadrupolar and octupolar wave modes. The fundamental framework governing these transitions is outlined as follows [44]:
@@
$\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)$
$=\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)+\Delta E_{s,s-1}\leavevmode\nobreak\ gr(rad)$ (15)
where $qu(rad)$ and $gr(rad)$ represents quantum radiation and gravitational radiation respectively. Also, note that throughout the article, ($qu$), ($gr$), and ($rad$) represent “quantum”, “gravitational”, and “radiation” respectively. We calculate the thermal quantum radiation using Eq. (7) as:
@@
and from Eq. (15), we get the gravitational radiation as:
$\Delta E_{s,s-1}\leavevmode\nobreak\ gr(rad)$
$=\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)-\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)$
@@
$=\frac{E_{p}}{2\sqrt{n_{s}}}\big{(}\frac{1}{4n_{s}}+\frac{1}{8n_{s}^{2}}+...\big{)}$ (17)
@@
$\Delta S_{s,s-1}\leavevmode\nobreak\ qu(rad)=2k_{B};$
$\Delta S_{s,s-1}\leavevmode\nobreak\ gr(rad)=2k_{B}\leavevmode\nobreak\ \big{(}\frac{1}{4n_{s}}+\frac{1}{8n_{s}^{2}}+...\big{)},$
@@
    Considering the straightforward physical expression of the acceleration $a_{s}$, linked to the dynamical process of a black hole’s evaporation using Eq.(10):
@@
    Using the previously stated equivalence $T_{p}=\frac{a_{p}}{ck_{B}}\hbar$, we can derive the remarkable expression that relates the thermal temperature and acceleration as:
    $T_{s}(rad)=\frac{E_{p}}{4k_{B}\sqrt{n_{s}}}$
@@
Note that, following the previous results, the basic relation (15) can be easily rewritten in the following forms:
$\Delta E_{s,s-1}(rad)=T_{s}\Delta S_{s,s-1};$
$\Delta E_{s,s-1}(rad)=F_{s}\Delta R_{s,s-1};$
$\Delta E_{s,s-1}(rad)=\frac{F_{s}}{2\pi}\Delta\lambda_{s,s-1};$
$\Delta E_{s,s-1}(rad)=p_{s}V_{s,s-1}.$ (23)
where $\Delta S_{s,s-1}$ is referred to (18), $\Delta R_{s,s-1}=2l_{p}(\sqrt{n_{s}}-\sqrt{n_{s-1}})$, $\Delta\lambda_{s,s-1}=2\pi\Delta R_{s,s-1}$ is referred to Compton’s wavelength of condensed light (1), $p_{s}=\frac{F_{s}}{A_{s}}$ is a PBH pressure, $\Delta V_{s,s-1}=A_{s}\Delta R_{s,s-1}$ is a quantum change of it’s volume.
All these unexpected equivalences indicate that the interplay between quantum and gravitational PBH radiation can be articulated through a diverse array of fundamental physical theories, including quantum mechanics, thermodynamics, hydrodynamics, and information theory, all of which incorporate a clearly defined gravitational aspect.
Moreover, from Eqs. (17) and (23), it is evident that this comprehensive physical description of the most important cosmic processes is governed by special form of the fundamental Heisenberg uncertainty principle:
$\Delta E_{s,s-1}\Delta t_{s}=\frac{E_{p}}{2\sqrt{n_{s}}}\big{(}1+\frac{1}{4n_{s}}+\frac{1}{8n_{s}^{2}}+...\big{)}2\pi t_{p}\sqrt{n_{s}}$
$=\frac{h}{2}\big{(}1+\frac{1}{4n_{s}}+\frac{1}{8n_{s}^{2}}+...\big{)}>\frac{h}{2}$ (24)
where $h$ is a non-reduced Planck constant.
The fundamental form encompasses gravitational components and is a direct consequence of the algebraic methodology imparted by Albert Einstein in his last lines [31].
@@ V PBHs Lifetime and Luminosity: Addressing the Information Loss Paradox
We now aim to determine the luminosity and lifetime of PBHs, in light of our newly developed formalism. By using the strong conditions of causality and spherical geometry one can readily calculate the transition time of PBHs from one quantum state to the underlying one. It is equal to half the length of the geodesic line divided by the speed of light [44],
$\Delta t_{s}=\frac{\pi R_{s}}{c}=2\pi t_{p}\sqrt{n_{s}}$ (25)
The aforementioned expression helps us to calculate the black hole lifetime, which can be readily found using the below-mentioned formula as:
$\tau_{s}=\sum_{n=0}^{n_{s}}2\pi t_{p}\sqrt{n}$
@@
where the usual lifetime and luminosity of the black hole can be expressed as: $\tau(Hawking)\simeq\frac{5120\pi G^{2}}{\hbar c^{4}}M^{3}$ and $L(Hawking)\simeq\frac{\hbar c^{6}}{15360\pi G^{2}M^{2}}$ [19, 51, 52]. Thus our model suggests that black holes undergoing quantum transitions differ substantially from Hawking’s semi-classical evaporation model. Quantum processes result in more intense emissions than continuous black-body radiation, reducing the black hole’s lifetime by a factor of 3840 and increasing its luminosity by the same factor due to discrete energy jumps. Consequently, without external energy input (e.g., through accretion) to offset this intense quantum emission, the black hole undergoes rapid evaporation. This finding challenges Hawking’s assumption that black holes radiate as black bodies under the Stefan-Boltzmann law, showing instead that quantum transitions produce significantly higher radiation intensities.
Now we know from Eq.(16) that the energy of each emitted particle is:
@@
Using Eq. (28), one can easily find that their emitted wavelength is much more than $R_{s}$ :
@@
Because the wavelength of the emitted particles is greater than the Schwarzschild radius, the PBH’s strong gravitational field does not prevent the particles from tunneling out. In our scenario, each quantum transition results in the emission of two particles, each containing a discrete amount of energy and information. As a result, each emission causes the PBH to lose two bits of quantum information. However, in conventional black hole evaporation, Hawking radiation is merely thermal and contains no information about the black hole’s internal state, resulting in the so-called information loss paradox. In contrast, our model may encode information in the emitted particles, allowing information about the black hole’s interior states to escape with each two-particle emission. As the PBH emits radiation, it loses mass, energy, and information in a quantized manner, resolving the information loss paradox by allowing information to leave the PBH in the form of emitted quantum bits (qubits). This can be easily deduced from the formula (19) as follows:
$I_{s}(rad)=\frac{\sum_{n=1}^{n_{s}}\Delta S_{s,s-1}qu(rad)}{k_{B}}=2n_{s}=I_{s}$ (30)
@@ VI Quantum Absorption of Free Light: Origins of Cosmic Disks and Jets
Considering the quantum characteristics of condensed light in PBHs, it is evident that only free light photons with a wavelength not exceeding the geodesic length of the PBHs are subject to accretion by quantum absorption. In other words, PBHs consume only free light or condensed light (such as from other black holes through collisions and mergers). All ordinary matter is melted in accretion disks and expelled in jets, accompanied by swirling magnetic fields through “funnel” and “slingshot” effects. As will be shown below, this is precisely what can explain the exceptionally early formation of the SMBHs from a huge reservoir of free light energy radiated by PBHs.
Referring to the previous section we conclude that the wavelength of an absorbed photon cannot exceed the geodesic length; thus we get a condition on wavelength:
$\lambda_{\gamma}\leavevmode\nobreak\ (absorp.)\leq 2\pi R_{s}=\lambda_{sm}$ (31)
@@
If, during the specified time interval $\Delta t_{s}$ (25), which corresponds to the duration of the quantum transition from energy level $n_{s}$ to $n_{s-1}$, the PBH does not absorb the aforementioned free light quantum of energy, it will continue to evaporate. If it receives, it grows in accordance with the law:
@@
It means when the temperature of the universe ($T_{\gamma}$) falls below twice the temperature of a PBH, it commences evaporation. Furthermore, it is essential to comprehend how, in the nascent universe, the demise of one generation of PBHs facilitated the emergence of the subsequent generation [3, 19, 51, 53]. We must adhere to these paramount cosmological criteria in our explanation of the formation of contemporary dark matter, which comprises black holes of asteroid-like mass [54]. Another significant consequence of these interactions is that an evaporating PBH can transfer its electromagnetic radiation energy to another PBH only if the mass of the latter is double that of the former [55],
@@
These correlations elucidate how, in the primordial phases of the universe’s growth, there occurred a cyclic transfer of energy from a dying, more abundant yet lighter generation of PBHs to a new, somewhat less abundant but heavier generation. This is essential for comprehending both the primordial formation of dark matter’s structure and the initial “clumping” of the universe, as these consistent cycles of dark matter generation coincided with the hyper-exponential proliferation of future SMBHs, which served as the centers for the formation of subsequent galaxies.
VII Conservation of Quantum Information and Gravitational Radiation from PBH merger
We find above that two-particle emission from the PBH accompanied by gravitational radiation can potentially resolve the ‘information loss paradox’. According to the definition provided by John Preskill [56] and Leonard Susskind [57], each emitted particle acts as a carrier of one bit of quantum information, consistent with the principles of Shannon entropy [58]. This information is encoded in two possible spin states, $|0>$ and $|1>$, as per the notation established by Paul Dirac [59]. This mechanism helps keep information safe and shows that the unitarity principle in quantum theory is conserved. This means that no information is lost during the evaporation of PBH.
In accordance with the law of information conservation for a binary black hole system ($BH_{1}$,$BH_{2}$), which is derived from the fundamental unitarity principle of quantum theory, we can write:
@@ VII Conservation of Quantum Information and Gravitational Radiation from PBH merger
From this, we get an interesting formula for the energy of gravitational radiation resulting from the collision and merging of a binary PBH system,
@@
This formulation, expressed in terms of quantum numbers, succinctly corroborates Hawking’s esteemed area theorem [60] concerning the colliding and merging of black holes. During the merger, the resultant black hole experiences a ”ringdown” phase, during which it emits gravitational waves while stabilizing into a final state. This process is similar to a struck bell that resonates and then diminishes into quiet, the black hole emits energy as it reaches equilibrium—thus the evocative name “ringdown.”
@@
Now, we obtain the general expression for total gravitational radiation energy for $N$ consecutive collisions and mergers of $N$ PBHs as :
@@
    When the mass of an SMBH at the center of a galaxy greatly exceeds the asteroid-like mass of PBHs situated in the SMBH’s “corona,” the dynamics of their mergers yield an interesting outcome. In these instances, nearly all the mass and energy of the tiny PBHs is converted into gravitational wave energy during the merger phase. This process demonstrates a significant transformation in which dark matter is permanently converted into dark energy within the universe. This process serves as a substantial validation of the famous last P. W. Anderson’s conjecture, that dark energy is the energy of gravitational radiation [62]. Moreover, as will be shown subsequently, these processes account for the universe’s present phase of accelerated expansion.
@@
    In contrast, Eqs. (40) and (41) suggest that the formation of an SMBH is unlikely during multiple collisions and mergers of PBHs with nearly equal masses, as most of their total energy would be dissipated as gravitational radiation. It can be inferred that SMBHs primarily originated from the significant light energy associated with radiative processes in the early universe. We will explain below that these critical insights are vital for precisely characterizing and comprehending the substantial role of PBHs and their gravitational radiation in cosmogenesis.
VIII Planck Scale to Expansion: The Quantum Foundations of Cosmology